Given g(x) = –2x + 1, find g(2a).
Answer by tutorcecilia(2152)

g(x) = –2x + 1
g(2a)=-2(2a)+1
g(2a)=-4a+1
.
checking:
-4a+1=-2x+1
-4a=-2x
-4a=-2(2a)
-4a=-4a [checks out]